OpenStudy (anonymous):

solve for the unknow for this equation. 8^z+1=32sqrt2

OpenStudy (anonymous):

write it properly

OpenStudy (anonymous):

\[8^{z+1}=32\sqrt{2}\]

OpenStudy (anonymous):

\[8^{z+1} = 32 \times \sqrt{2}\] \[2^{3 \times (z+1)} = 2^{5} \times 2^{\frac{1}{2}}\] \[2^{ 3z+3} = 2^{5+ \frac{1}{2}} \] Both sides base is common so \[{ 3z+3} = {5+ \frac{1}{2}} \] 6z + 6 = 10 + 1 6z = 11-6 \[z = \frac{5}{6} \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

let me process how you did this

OpenStudy (anonymous):

is there a formula I need to understand for this?